OpenStudy (anonymous):

domain of 4/sqrt(x+3)

OpenStudy (luigi0210):

Hello @asdf1234asdf Welcome to openstudy :) To find the domain, first set that x+3 equal to 0

OpenStudy (anonymous):

@Luigi0210 so 4sqrt3/3?

OpenStudy (anonymous):

solve \[x+3>0\] for \(x\) in exactly one step

OpenStudy (anonymous):

you can't take the square root of a negative number you can't divide by zero therefor \(x+3\) must be greater than 0

OpenStudy (anonymous):

@satellite73 so the 4 has nothing to do with finding the domain?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

would be the same if you have \[\frac{7}{\sqrt{x+3}}\] or \[\frac{\pi}{\sqrt{x+3}}\]

OpenStudy (anonymous):

@satellite73 would i use x+3>0 for interval notation? or would i need greater than or equal to?

OpenStudy (anonymous):

you would solve \(x+3>0\) for \(x\)